OpenStudy (anonymous):

Determine the solution set of (x-3)^2=49

OpenStudy (anonymous):

Take the square root of both sides to eliminate the square \[x-3= \pm 7\]

OpenStudy (anonymous):

So now you have two simple equations to solve x-3=7 and x-3=-7